Suppose we observe the following rates: 1R1 = 10%, 1R2 = 14%, and E(2r1) = 18%. If the liquidity premium theory of the term structure of interest rates holds, what is the liquidity premium for year 2? (Round your intermediate calculations to 5 decimal places and final answer to 3 decimal places. (e.g., 32.161))
Answer #1

1 + 1R2= {(1 + 1R1)(1 + E(2r1) + L2)}1/2

1.14 = (1.10)(1.18 + L2)1/2

(1.14)2/1.10 = 1.18 + L2

1.29960/1.1 = 1.18 + L2

L2 = 0.145%